If the numerator of a fraction is 8 less than the denominator of the fraction and the value of the fraction is 3/5, what's the numerator of this fraction?
A)9
B)12
C)15
D)16
E)20

Answer by Boreal(15235)

x=numerator
x+8 is denominator
x/(x+8)=3/5
3x+24=5x, by cross-multiplication
2x=24
x=12, the numerator.
12/20 is the fraction.